Two circles have the same center but one circle has a radius 5 cm greater than the other. The area of the inner circle is 228 cm2 less than the outer circle. What is the radius of the smaller circle in centimeters?

Answer by Boreal(15235)

Radius of smaller is x
of larger is x+5
area of smaller/inner is pi*x^2
area of larger is pi(x^2+10x+25)
smaller + 228 cm^2=larger
so pi(x^2)+228=pix^2+pi(10x+pi*25)
228=pi(10x+25)
72.57=10x+25
10x=47.57
x=4.76 cm radius smaller circle with area 71.18 cm^2 ANSWER
x+5=9.76 cm radius larger circle with area 299.26 cm^2, 228 cm^2 larger given rounding.